OpenStudy (anonymous):

How to show that for any vector in R3, (v dot (Av) is greater than or equal to 0

OpenStudy (anonymous):

sorry I don't know how to do this

OpenStudy (anonymous):

do you know linear algebra? I have a super hard take home test im trying to do

OpenStudy (anonymous):

give me a question and ill see if I can do it

OpenStudy (anonymous):

have you taken a linear algebra course? i'm typing one right now

OpenStudy (anonymous):

Let A and B be nxn maricies. Suppose v is an eigenvector of A associated with the eigenvalue x and is also an eigenvector of B associated with the eigenvalue y. Show that x+y is an eigenvalue of A+B

OpenStudy (anonymous):

nope never seen anything like it sorry

OpenStudy (anonymous):

yess, currently in linear algebra, haven't read the question yet. Lemme see

OpenStudy (anonymous):

Are you sure about that problem statement? Because it isnt true. As a matter of fact take v = (1,1,1) and\[A = \left[\begin{matrix}-1 &0 & 0 \\ 0 & -1 & 0\\ 0 & 0 & -1\end{matrix}\right]\]

OpenStudy (anonymous):

unless, of course, you're referring to an eigenvector by using v, but then it wouldn't be any vector in R3

OpenStudy (anonymous):

Because v is an eigenvector of A associated with the eigenvalue x, then : \[Av=xv\] and because v is an eigenvector of B associated with the eigenvalue y, then : \[Bv=yv\] Hence : \[(A+B)v=Av+Bv=xv+yv=(x+y)v\] Therefor x+y is an eigenvalue of A+B